A particle of mass m moving in the x-direction with speed 2v is hit by another particle of mass 2m moving in the y-direction with speed v. If the collision is perfectly inelastic, the percentage loss in the energy during the collision is close to : 

Grade:12

1 Answers

Adarsh
6 years ago
initial energy of system before collision = ½.m.4v2+ ½.2m.v2 =  3.m.v2
in inelastic collision two particle will stick together and their momentum will be conserved .
so,  4mv = 3m.v” => v” = 4/3v.
so the velocity after collision is 4/3 times the velocity before collision.
now, energy after collision = ½.3m.(4/3v)2= 8m/3.v2
%loss in kinetic energy = 100 – 3.m.v2/8/3.m.v2 x 100 = 11.11%
 
Thanks and Regards 
Adarsh